OpenStudy (anonymous):

g(x)=square root of 5x-4 h(x)4x^2+7 Find: h(g(1))

OpenStudy (espex):

Start by substituting your '1' into your inner function, then take the output and insert it into your outer function.

OpenStudy (anonymous):

so do square root of 5 and then subtract 4 first?

OpenStudy (espex):

You have \[\sqrt{5x-4}\] and you are told that x=1, so start with that math, what do you get for an answer?

OpenStudy (anonymous):

i dont have a calculator so would it be .5 or 1?

OpenStudy (espex):

Well you have 5*1-4=5-4=1, then you have a sqrt of 1, which is 1. Take this answer and plug it in for the 'x' of your h function.